# Grounding DINO: Marrying DINO with Grounded Pre-Training for Open-Set Object Detection

### v1

@@ Abstract
In this paper, we present an open-set object detector, called Grounding DINO, by marrying Transformer-based detector DINO with grounded pre-training, which can detect arbitrary objects with human inputs such as category names or referring expressions. The key solution of open-set object detection is introducing language to a closed-set detector for open-set concept generalization. To effectively fuse language and vision modalities, we conceptually divide a closed-set detector into three phases and propose a tight fusion solution, which includes a feature enhancer, a language-guided query selection, and a cross-modality decoder for cross-modality fusion. While previous works mainly evaluate open-set object detection on novel categories, we propose to also perform evaluations on referring expression comprehension for objects specified with attributes. Grounding DINO performs remarkably well on all three settings, including benchmarks on COCO, LVIS, ODinW, and RefCOCO/+/g. Grounding DINO achieves a $52.5$ AP on the COCO detection zero-shot transfer benchmark, i.e., without any training data from COCO. It sets a new record on the ODinW zero-shot benchmark with a mean $26.1$ AP. Code will be available at https://github.com/IDEA-Research/GroundingDINO.
@@ 3.3 Cross-Modality Decoder
image_features:Tensor.shape:(bs,num_img_tokens,ndim)
text_features:Tensor.shape:(bs,num_text_tokens,ndim)
@@
topk_proposals_idx:Tensor.shape:(bs,num_query)
@@ 3.5 Loss Function
[…]
@@
Grounding DINO T Swin-T O365 46.7 56.9 / -
Grounding DINO T Swin-T O365,GoldG 48.1 57.1 / -
Grounding DINO T Swin-T O365,GoldG,Cap4M 48.4 57.2 / -
Grounding DINO L Swin-L O365,OI[19],GoldG 52.5 62.6 / 62.7 (63.0 / 63.0)*
Grounding DINO L Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 60.7 62.6 / -
@@
Grounding DINO T Swin-T O365,GoldG 25.6 14.4/19.6/32.2
Grounding DINO T Swin-T O365,GoldG,Cap4M 27.4 18.1/23.3/32.7
Grounding DINO L Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 33.9 22.2/30.7/38.8
@@ 4.1 Setup
Implementation Details We trained two model variants, Grounding DINO T with Swin-T [32] as an image backbone, and Grounding DINO L with Swin-L [32] as an image backbone. We leveraged BERT-base [8] from Hugging Face [51] as text backbones. As we care more about the model performance on novel classes, we list zero-shot transfer and referring detection results in the main text. More implementation details are available in the Appendix Sec. A.
@@ 4.2.2 LVIS Benchmark
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 20.0 9.5
Grounding DINO T(Ours) Swin-T 172M O365,GoldG,Cap4M 22.3 11.9
@@
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 46.4 51.1
@@
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 70.7 76.2
4.2.3 ODinW Benchmark
ODinW (Object Detection in the Wild) [23] is a more challenging benchmark to test model performance under real-world scenarios. It collects more than $35$ datasets for evaluation. We report three settings, zero-shot, few-shot, and full-shot results in Table 4. Grounding DINO performs well on this benchmark. With only O365 and GoldG for pre-train, Grounding DINO T outperforms DINO on few-shot and full-shot settings. Impressively, Grounding DINO with a Swin-T backbone outperforms DINO with Swin-L on the full-shot setting. Grounding DINO outperforms GLIP under the same backbone for the zero-shot setting, comparable with GLIPv2 [59] without any new techniques like masked training. The results show the superiority of our proposed models. Grounding DINO L set a new record on ODinW zero-shot with a $26.1$ AP, even outperforming the giant Florence models [55]. The results show the generalization and scalability of Grounding DINO.
@@ 4.2.3 ODinW Benchmark
Grounding DINO T (Ours) Swin-T O365,GoldG 50.41 57.24 43.21 51.40 57.59 45.81 67.46 67.13
Grounding DINO T (Ours) Swin-T O365,GoldG,RefC 73.98 74.88 59.29 66.81 69.91 56.09 71.06 72.07
Grounding DINO T (Ours) Swin-T O365,GoldG,RefC 89.19 91.86 85.99 81.09 87.40 74.71 84.15 84.94
Grounding DINO L (Ours)* Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 90.56 93.19 88.24 82.75 88.95 75.92 86.13 87.02
@@ 4.5 Transfer from DINO to Grounding DINO
Recent work has presented many large-scale image models for detection with DINO architecture [^6] See https://github.com/IDEA-Research/detrex for some model instances. It is computationally expensive to train a Grounding DINO model from scratch. However, the cost can be significantly reduced if we leverage pre-trained DINO weights. Hence, we conduct some experiments to transfer pre-trained DINO to Grounding DINO models. We freeze the modules co-existing in DINO and Grounding DINO and fine-tune the other parameters only. (We compare DINO and Grounding DINO in Sec. E.) The results are available in Table 7.
It shows that we can achieve similar performances with Grounding DINO training only text and fusion blocks using a pre-trained DINO. Interestingly, the DINO-pre-trained Grounding DINO outperforms standard Grounding DINO on LVIS under the same setting. The results show that there might be much room for model training improvement, which will be our future work to explore.
With a pre-trained DINO initialization, the model converges faster than Grounding DINO from scratch, as shown in Fig. 5.
@@ 6 Acknowledgement
We thank the authors of GLIP [26]: Liunian Harold Li, Pengchuan Zhang, and Haotian Zhang for their helpful discussions and instructions. We also thank Tiancheng Zhao, the author of OmDet [61], and Jianhua Han, the author of DetCLIP [53], for their kind replies for their model details.
@@ Appendix B Data Usage
There are two versions of the O365 dataset, which we termed O365v1 and O365v2, respectively. O365v1 is a subset of O365v2. O365v1 contains about 600K images, while O365v2 contains about 1.7M images. Following previous works [26, 53], we pre-train the Grounding DINO T on O365v1 for a fair comparison. The Grounding DINO L is pre-trained on O365v2 for a better result.
@@ Appendix G Marry Grounding DINO with Stable Diffusion
Grounding DINO L finishes the detection task in Fig. 1. The results in Fig. 1 (b) are generated by two processes. First, we detect objects with Grounding DINO. Then we generate masks by masking out the detected objects or backgrounds. After that, we feed original images, image masks, and generation prompts to the stable diffusion models [42] to render new images. We use the releases checkpoints in https://github.com/Stability-AI/stablediffusion for new image generation.
@@ Appendix I Model Efficiency
We compare the model size and efficiency between Grounding DINO T and GLIP-T in Table 14. The results show that our model has a smaller parameter size and better efficiency than GLIP.

### v3

@@ Abstract
In this paper, we present an open-set object detector, called Grounding DINO, by marrying Transformer-based detector DINO with grounded pre-training, which can detect arbitrary objects with human inputs such as category names or referring expressions. The key solution of open-set object detection is introducing language to a closed-set detector for open-set concept generalization. To effectively fuse language and vision modalities, we conceptually divide a closed-set detector into three phases and propose a tight fusion solution, which includes a feature enhancer, a language-guided query selection, and a cross-modality decoder for cross-modality fusion. While previous works mainly evaluate open-set object detection on novel categories, we propose to also perform evaluations on referring expression comprehension for objects specified with attributes. Grounding DINO performs remarkably well on all three settings, including benchmarks on COCO, LVIS, ODinW, and RefCOCO/+/g. Grounding DINO achieves a $52.5$ AP on the COCO detection zero-shot transfer benchmark, i.e., without any training data from COCO. After fine-tuning with COCO data, Grounding DINO reaches $63.0$ AP. It sets a new record on the ODinW zero-shot benchmark with a mean $26.1$ AP. Code will be available at https://github.com/IDEA-Research/GroundingDINO.
@@ 3.3 Cross-Modality Decoder
image_features:(bs,num_img_tokens,ndim)
text_features:(bs,num_text_tokens,ndim)
@@
topk_proposals_idx:(bs,num_query)
@@ 3.5 Loss Function
Faster R-CNN RN50-FPN - - 40.2 / -
Faster R-CNN RN101-FPN - - 42.0 / -
DyHead-T [5] Swin-T - - 49.7 / -
DyHead-L [5] Swin-L - - 58.4 / 58.7
DyHead-L [5] Swin-L O365,ImageNet21K - 60.3 / 60.6
SoftTeacher [52] Swin-L O365,$S$-COCO - 60.7 / 61.3
DINO(Swin-L) [58] Swin-L O365 - 62.5 / -
@@
Grounding-DINO-T (Ours) Swin-T O365 46.7 56.9 / -
Grounding-DINO-T (Ours) Swin-T O365,GoldG 48.1 57.1 / -
Grounding-DINO-T (Ours) Swin-T O365,GoldG,Cap4M 48.4 57.2 / -
Grounding-DINO-L (Ours) Swin-L O365,OI[19],GoldG 52.5 62.6 / 62.7 (63.0 / 63.0)*
Grounding-DINO-L (Ours) Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 60.7 62.6 / -
@@
Grounding-DINO-T Swin-T O365,GoldG 25.6 14.4/19.6/32.2
Grounding-DINO-T Swin-T O365,GoldG,Cap4M 27.4 18.1/23.3/32.7
Grounding-DINO-L Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 33.9 22.2/30.7/38.8
@@ 4.1 Setup
Implementation Details We trained two model variants, Grounding-DINO-T with Swin-T [32], and Grounding-DINO-L with Swin-L [32] as an image backbone, respectively. We leveraged BERT-base [8] from Hugging Face [51] as text backbones. As we focus more on the model performance on novel classes, we list zero-shot transfer and referring detection results in the main text. More implementation details are available in the Appendix Sec. A.
@@ 4.2 Zero-Shot Transfer of Grounding DINO
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 20.0 9.5
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG,Cap4M 22.3 11.9
@@
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 46.4 51.1
@@
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 70.7 76.2
ODinW Benchmark
ODinW (Object Detection in the Wild) [23] is a more challenging benchmark to test model performance under real-world scenarios. It collects more than $35$ datasets for evaluation. We report three settings, zero-shot, few-shot, and full-shot results in Table 4. Grounding DINO performs well on this benchmark. With only O365 and GoldG for pre-train, Grounding-DINO-T outperforms DINO on few-shot and full-shot settings. Impressively, Grounding DINO with a Swin-T backbone outperforms DINO with Swin-L on the full-shot setting. Grounding DINO outperforms GLIP under the same backbone for the zero-shot setting, comparable with GLIPv2 [59] without any new techniques like masked training. The results show the superiority of our proposed models. Grounding-DINO-L set a new record on ODinW zero-shot with a $26.1$ AP, even outperforming the giant Florence models [55]. The results show the generalization and scalability of Grounding DINO.
@@
Grounding-DINO-T (Ours) Swin-T O365,GoldG 50.41 57.24 43.21 51.40 57.59 45.81 67.46 67.13
Grounding-DINO-T (Ours) Swin-T O365,GoldG,RefC 73.98 74.88 59.29 66.81 69.91 56.09 71.06 72.07
Grounding-DINO-T (Ours) Swin-T O365,GoldG,RefC 89.19 91.86 85.99 81.09 87.40 74.71 84.15 84.94
Grounding-DINO-L (Ours)* Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 90.56 93.19 88.24 82.75 88.95 75.92 86.13 87.02
@@ 4.5 Transfer from DINO to Grounding DINO
Recent work has presented many large-scale image models for detection with DINO architecture [^6] See model instances at https://github.com/IDEA-Research/detrex . It is computationally expensive to train a Grounding DINO model from scratch. However, the cost can be significantly reduced if we leverage pre-trained DINO weights. Hence, we conduct some experiments to transfer pre-trained DINO to Grounding DINO models. We freeze the modules co-existing in DINO and Grounding DINO and fine-tune the other parameters only. (We compare DINO and Grounding DINO in Sec. E.) The results are available in Table 7.
It shows that we can achieve similar performances with Grounding-DINO-Training only text and fusion blocks using a pre-trained DINO. Interestingly, the DINO-pre-trained Grounding DINO outperforms standard Grounding DINO on LVIS under the same setting. The results show that there might be much room for model training improvement, which will be our future work to explore. With a pre-trained DINO initialization, the model converges faster than Grounding DINO from scratch, as shown in Fig. 5. Notably, we use the results without exponential moving average (EMA) for the curves in Fig. 5, which results in a different final performance that in Table 7. As the model trained from scratch need more training time, we only show results of early epochs.
@@ 6 Acknowledgement
We thank the authors of GLIP [26]: Liunian Harold Li, Pengchuan Zhang, and Haotian Zhang for their helpful discussions and instructions. We also thank Tiancheng Zhao, the author of OmDet [61], and Jianhua Han, the author of DetCLIP [53], for their response on their model details. We thank He Cao of The Hong Kong University of Science and Technology for his helps on diffusion models.
@@ Appendix B Data Usage
There are two versions of the O365 dataset, which we termed O365v1 and O365v2, respectively. O365v1 is a subset of O365v2. O365v1 contains about 600K images, while O365v2 contains about 1.7M images. Following previous works [26, 53], we pre-train the Grounding-DINO-T on O365v1 for a fair comparison. The Grounding-DINO-L is pre-trained on O365v2 for a better result.
@@ Appendix G Marry Grounding DINO with Stable Diffusion
Grounding DINO L finishes the detection task in Fig. 1. The results in Fig. 1 (b) are generated by two processes. First, we detect objects with Grounding DINO. Then we generate masks by masking out the detected objects or backgrounds. After that, we feed original images, image masks, and generation prompts to the stable diffusion models [42] to render new images. We use the releases checkpoints in https://github.com/Stability-AI/stablediffusion for new image generation. More results are available in Figure 8.
Using GLIGEN for Grounded Generation
To enable fine-grained image editing, we combine the Grounding DINO with GLIGEN [GLIGEN]. GLIGEN supports grounding results as inputs and can generate objects on specific positions. We can assign each bounding box an object with GLIGEN, as the cake and dog in Figure 9 (c) (d). Moreover, GLIGEN can full fill each bounding box, which results in better visualization, as that in Figure 9 (a) (b). For example, we use the same generative prompt in Figure 8 (b) and Figure 9 (b). The GLIGEN results ensure each bounding box with an object and fulfills the detected regions.
Figure 8: Combination of Grounding DINO and Stable Diffusion. We first detect objects with Grounding DINO and then perform image inpainting with Stable Diffusion. * The input human face in the row (e) is generated by StyleGAN.
Figure 9: Combination of Grounding DINO and GLIGEN. We first detect objects with Grounding DINO and then perform image inpainting with GLIGEN. * The phrase prompts are separated by semicolons. We assign phrase prompts to bounding boxes randomly.
@@ Appendix I Model Efficiency
We compare the model size and efficiency between Grounding-DINO-T and GLIP-T in Table 14. The results show that our model has a smaller parameter size and better efficiency than GLIP.